OpenStudy (anonymous):

what is the area of a circle with the diameter of 3.8

OpenStudy (e.mccormick):

Area of a circle is \(\pi r^2\). Do you know how to get the r from what you were given?

OpenStudy (anonymous):

thnkz

OpenStudy (anonymous):

Area=pi(3.14)*(1.9)^2=11.34

OpenStudy (anonymous):

what makes a triangle a right triangle ?

OpenStudy (e.mccormick):

Do you know what a right angle is?

OpenStudy (anonymous):

yes , a triangle with 90 degree angle!!!

OpenStudy (e.mccormick):

There you go!